# Kernelized Decoded Quantum Interferometry

### v1

@@ C.2 Justification via discrete stationary phase
The PPC property can be motivated by discrete stationary-phase and linear canonical transform identities. Quadratic-phase kernels exactly diagonalize quadratic exponential sums and approximately diagonalize higher-degree polynomial phases, with tails controlled by Weyl-type exponential-sum bounds; see, for example, [HealyLCT2016, Rabiner1969CZT] for LCT/fractional Fourier transform identities and [JordanNature2025, JordanArxiv2024] for how the OPI shaping map inherits low-degree algebra from the underlying polynomial $h$. A detailed coding of these classical estimates into the DQI context is beyond our scope; Definition C.1 encapsulates the resulting concentration behavior in a form suitable for our head-mass analysis.
@@ D.1 Setup and notation
Let $B\in\{0,1\}^{m\times n}$ be a sparse parity-check matrix with left/right degree distributions $(\lambda,\rho)$ and girth at least $\Omega(\log n)$ on typical instances. The DQI reduction maps the shaped state $g(x)=P(f(x))$ through the interferometer $\mathcal{F}$ to amplitudes supported on rows of $B$ and local combinations thereof; see [JordanNature2025, JordanArxiv2024] for details. We partition the $n$ qubits into $N/b$ disjoint blocks of width $b=O(1)$ (for clarity we assume $b\mid N$), and we consider block-local kernels of the form
@@ Appendix E Smoothness of the Optimization Landscape
Let $\Sigma_{K}(\theta)=\braket{\psi_{out}|\Pi_{Head}|\psi_{out}}$ be the head mass probability. The magnitude of the gradient with respect to $\theta$ is bounded by:

### v2

@@ C.2 Justification via discrete stationary phase
The PPC property can be motivated by discrete stationary-phase and linear canonical transform identities. Quadratic-phase kernels exactly diagonalize quadratic exponential sums and approximately diagonalize higher-degree polynomial phases, with tails controlled by Weyl-type exponential-sum bounds; see, for example, [HealyLCT2016, Rabiner1969CZT] for LCT/fractional Fourier transform identities and [JordanNature2025] for how the OPI shaping map inherits low-degree algebra from the underlying polynomial $h$. A detailed coding of these classical estimates into the DQI context is beyond our scope; Definition C.1 encapsulates the resulting concentration behavior in a form suitable for our head-mass analysis.
Lemma C.2 (PPC for affine OPI instances)
Let $p$ be prime and consider an OPI instance over $\mathbb{F}_{p}$ with evaluation points $\{0,1,\dots,p-1\}$ and a polynomial $f(x)=ax+b$ of degree at most $1$. Suppose the shaping map of Section IV is chosen so that the shaped amplitudes before the kernel are proportional to
$\psi_{x}\propto e^{2\pi if(x)/p},\qquad x\in\mathbb{F}_{p}.$ (45)
Then with the identity kernel $K=\mathbb{I}$ and the pp–ary QFT $F_{p}$, the post–interference spectrum $F_{p}\psi$ is supported on a single mode. In particular, Definition C.1 holds with degree bound $d=1$, head size $k=1$, and tail parameter $\varepsilon_{p}(1,1)=0$.
proof C.2
A direct computation shows that
$(F_{p}\psi)_{m}\;\propto\;\sum_{x\in\mathbb{F}_{p}}e^{2\pi i(ax+b)/p}e^{-2\pi imx/p}\;=\;e^{2\pi ib/p}\sum_{x\in\mathbb{F}_{p}}e^{2\pi i(a-m)x/p}.$ (46)
The inner sum vanishes unless $m\equiv a\pmod{p}$, in which case it equals $p$. Thus all spectral mass is concentrated on the unique index $m^{\star}$ with $m^{\star}\equiv a\pmod{p}$, and the normalized spectrum satisfies $|(F_{p}\psi)_{m^{\star}}|^{2}=1$ and $(F_{p}\psi)_{m}=0$ for $m\neq m^{\star}$. This is exactly the PPC property with $k=1$ and zero tail.
@@ D.1 Setup and notation
Let $B\in\{0,1\}^{m\times n}$ be a sparse parity-check matrix with left/right degree distributions $(\lambda,\rho)$ and girth at least $\Omega(\log n)$ on typical instances. The DQI reduction maps the shaped state $g(x)=P(f(x))$ through the interferometer $\mathcal{F}$ to amplitudes supported on rows of $B$ and local combinations thereof; see [JordanNature2025] for details. We partition the $n$ qubits into $N/b$ disjoint blocks of width $b=O(1)$ (for clarity we assume $b\mid N$), and we consider block-local kernels of the form
@@ D.3 From BLA to weighted head mass
4. Specialization to regular $(3,6)$ LDPC ensembles on the BEC
For regular LDPC ensembles on the binary erasure channel (BEC), the effect of block-local alignment can be expressed directly at the level of the density-evolution (DE) recursion. We record this specialization for the canonical $(d_{\ell},d_{r})=(3,6)$ family used in the main text.
Recall that for a regular $(3,6)$ ensemble on $\mathrm{BEC}(\varepsilon)$, the DE map has the explicit form
$\phi_{\varepsilon}(x)\;=\;\varepsilon\,\lambda\bigl(1-\rho(1-x)\bigr),\qquad\lambda(z)=z^{2},\;\;\rho(z)=z^{5},$ (62)
so that
$\phi_{\varepsilon}(x)=\varepsilon\Bigl(1-(1-x)^{5}\Bigr)^{2}.$
Let $\varepsilon^{\star}_{\mathrm{BEC}}\approx 0.429$ denote the asymptotic BP threshold for this ensemble, characterized by the usual tangency condition: there exists a point $x^{\star}\in(0,1)$ such that
$\phi_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})=x^{\star},\qquad\phi^{\prime}_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})=1.$ (63)
In the k–DQI setting we model the effect of BLA on the DE surrogate by an effective erasure parameter
$\varepsilon_{\mathrm{eff}}\;=\;\varepsilon\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr),\qquad 0<\kappa_{3,6}\leq 1,$ (64)
where $\Delta\Sigma_{\mathrm{loc}}>0$ is the additive head-mass gain per block from Definition D.1 and $\kappa_{3,6}$ is a calibration constant that depends only on the $(3,6)$ ensemble and the chosen block width $b$. We implicitly assume $\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}<1$ so that the effective erasure parameter remains in the physically meaningful regime. Numerically, $\kappa_{3,6}$ is obtained by matching the k–DQI head-mass gain to the shift of the DE waterfall curve (see Fig. 2 in the main text). Mathematically, the key property is simply that $\varepsilon_{\mathrm{eff}}$ is strictly smaller than $\varepsilon$ and depends linearly on $\Delta\Sigma_{\mathrm{loc}}$.
Proposition D.1 (BLA-induced DE improvement for regular $(3,6)$ ensembles)
Consider the regular $(3,6)$ ensemble on $\mathrm{BEC}(\varepsilon)$ with DE map $\phi_{\varepsilon}$ as in (62). Assume that the block-local alignment property of Definition D.1 holds with gain $\Delta\Sigma_{\mathrm{loc}}>0$ and that the corresponding DE surrogate uses the effective erasure parameter $\varepsilon_{\mathrm{eff}}$ from (64). Let $\varepsilon=\varepsilon^{\star}_{\mathrm{BEC}}$ be the BP threshold satisfying (63). Then for any $\Delta\Sigma_{\mathrm{loc}}>0$ we have
$\phi_{\varepsilon_{\mathrm{eff}}}(x^{\star})$ $=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\phi_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})\;=\;\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)x^{\star}\;<\;x^{\star},$ (65)
$\phi^{\prime}_{\varepsilon_{\mathrm{eff}}}(x^{\star})$ $=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\phi^{\prime}_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})\;=\;\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\;<\;1.$ (66)
In particular, the DE recursion with parameter $\varepsilon_{\mathrm{eff}}$ is strictly contractive in a neighbourhood of $x^{\star}$ and thus lies in the BP-success regime. The distance below the diagonal and the margin $1-\phi^{\prime}_{\varepsilon_{\mathrm{eff}}}(x^{\star})$ grow linearly with $\Delta\Sigma_{\mathrm{loc}}$.
proof D.2
The DE map (62) is linear in $\varepsilon$, i.e., $\phi_{\varepsilon_{\mathrm{eff}}}(x)=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\phi_{\varepsilon}(x)$ for every $x\in[0,1]$ whenever (64) holds. The same linear relation holds for the derivative. Specializing to $\varepsilon=\varepsilon^{\star}_{\mathrm{BEC}}$ and $x=x^{\star}$ and using the tangency conditions (63) yields
$\phi_{\varepsilon_{\mathrm{eff}}}(x^{\star})$ $=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\phi_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)x^{\star},$
$\phi^{\prime}_{\varepsilon_{\mathrm{eff}}}(x^{\star})$ $=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr)\phi^{\prime}_{\varepsilon^{\star}_{\mathrm{BEC}}}(x^{\star})=\bigl(1-\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\bigr).$
Since $0<\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\leq 1$ by assumption, both quantities are strictly smaller than their threshold counterparts, proving (65)–(66). The fact that the recursion is contractive in a neighbourhood of $x^{\star}$ follows from standard one-dimensional fixed-point theory: a differentiable map with slope strictly less than 1 in modulus is locally attracting. The linear dependence on $\Delta\Sigma_{\mathrm{loc}}$ is explicit in the prefactors.
Remark D.1 (Calibration of $\kappa_{3,6}$ and connection to simulations)
In practice, $\kappa_{3,6}$ is estimated once by matching the head-mass gain $\Delta\Sigma_{\mathrm{loc}}$ to the observed shift of the DE waterfall curve in numerical experiments (see Fig. 2). Proposition D.1 then shows that any positive calibrated value of $\kappa_{3,6}$ suffices to move the DE map strictly below the diagonal at the original threshold point and to dampen the derivative peak. Thus, within the DE surrogate, the BLA-induced gain $\Delta\Sigma_{\mathrm{loc}}$ translates directly into a provable margin of BP stability.
@@ Appendix E Smoothness of the Optimization Landscape
Let $\Sigma_{K}(\theta)=\braket{\Phi(\theta)|\Pi_{Head}|\Phi(\theta)}$ be the head mass probability. The magnitude of the gradient with respect to $\theta$ is bounded by: